OpenStudy (fredoneedshelp):

Solve for M F=mv2/r this is a fraction btw and the 2 is an exponent

OpenStudy (iambatman):

\[F= \frac{ mv^2 }{ r }\] ah good ole centripetal force, alright any idea how to begin?

OpenStudy (fredoneedshelp):

no

OpenStudy (fredoneedshelp):

i need help on these

OpenStudy (iambatman):

Ok so what happens if you multiply both sides by r?

OpenStudy (fredoneedshelp):

it turns to r?

OpenStudy (fredoneedshelp):

???

OpenStudy (fredoneedshelp):

is it mv^2?

OpenStudy (iambatman):

\[\color{red}r F = \frac{ mv^2 }{ r } \color{red}{r}\] what happens to the r on the right side

OpenStudy (fredoneedshelp):

uhh

OpenStudy (fredoneedshelp):

turns to rf/v^2

OpenStudy (iambatman):

Notice the r gets cancelled on the right, the idea is to get "m" on its own so when we do something to one side we have to do it to the other, or if you do it to the left side you must also do it to the right, then we're left with \[F r = mv^2\] now we have that v^2 what can we do?

OpenStudy (fredoneedshelp):

fr/v^2

OpenStudy (iambatman):

Yes!

OpenStudy (fredoneedshelp):

thats the answer?

OpenStudy (iambatman):

Sounds good!

OpenStudy (iambatman):

\[m = \frac{ Fr }{ v^2 }\]

OpenStudy (fredoneedshelp):

I got the answer right thanks